A current of 5.75 A flows in the coil of wire from an applied voltage of 120 V. The motor operating at normal speed creates a back EMF of 32.0 V. What is the resistance of the coil?

1
Expert's answer
2021-12-08T09:43:24-0500

"\u03f5i=V_S\u2212IR"


"R=-[\frac{ϵi-V_S}{I}]"

Where;

ϵi=back EMF=32V

"V_s" ==applied voltage=120v

I= current=5.75A

R=resistance


"R=-[\frac{32-120}{5.75}]=15.3ohms"

R=15.3 ohms